Circular No. 8960

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(136472) MAKEMAKE The IAU’s Committee on Small Body Nomenclature and Working

Group for Planetary System Nomenclature have agreed to give to the transneptunian dwarf planet, or ‘plutoid’, (136472) 2005 FY_{9} (cf. IAUC 8577) the name “Makemake”, as announced on MPC 63395.

2008 KV_{42}

An asteroidal object (mag r = 23.7–23.8) discovered by B. Gladman, J. Kavelaars, and J.-M. Petit on CCD images obtained on May 31 with the 3.5-m Canada-France-Hawaii telescope (and confirmed via observations made at Cerro Tololo and Mt. Hopkins through July 8) has a retrograde orbit (i = 103^{o}) with q ∼ 20 AU, Q ∼ 72 AU, and T ∼ 2034 (currently r = 32 AU). Details are given on MPEC 2008-O02.

COMETS C/2008 H2–H5 (SOHO) Further to IAUC 8955, additional apparent comets have been found

on SOHO website images; all are Kreutz sungrazers except for C/2008 H4 (Meyer group), which was small and elongated but not diffuse (mag ∼ 7.5). C/2008 H2 was stellar in appearance and very faint (mag ∼ 8) but very long-lasting in the SOHO images. C/2008 H3 was teardrop-shaped and quite bright (mag ∼ 6) in C2 images. C/2008 H5 was small with a short tail (mag ∼ 6.5) in C2 images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 H2 Apr. 16.588 1^{h}47.^{m}8 + 9^{o}26′ C2  HS 2008-L29
C/2008 H3  16.721 1 52.0 + 9 16 C3/2 RM 2008-L29
C/2008 H4 18.354 1 51.7 +12 32 C2 RK 2008-L47
C/2008 H5  25.513 2 23.3 +11 56 C3/2 HS 2008-L47

COMET P/2008 L2 (HILL) The following improved orbital elements from MPC 63377 show this

comet (cf. IAUC 8953) to be of short period.

T = 2008 Aug. 17.7902 TT  ω = 140.^{o}9570
e = 0.621446  Ω = 218.1044 }2000.0
q = 2.321935 AU  i = 25.8459
a = 6.133699 AU n^{o} = 0.0648814 P = 15.2 years
2008 July 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green